# Relative monodromy of ramified sections on abelian schemes

### v3

@@ 2.3 Monodromy of periods
Recall that the monodromy group of the universal family $\textrm{Mon}(\mathfrak{A}_{g})$ is a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$. When the modular map $p:S\to\mathbb{A}_{g}$ is finite, the monodromy group $\textrm{Mon}(\mathcal{A})$ of $\mathcal{A}\to S$ is still a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$ thanks to Equation 10; in particular, $\textrm{Mon}(\mathcal{A})$ is Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$. Moreover, when the image of the modular map $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$, the monodromy group $\textrm{Mon}(\mathcal{A})$ is still Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$ by [3, Remark 10.1.2].
@@
Assume that there exists a non-trivial subvector space $W$ of $\mathbb{Q}^{2g}$ which is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Z})$. Since the symplectic group $\textrm{Sp}_{2g}$ is a simple algebraic Lie group defined over $\mathbb{Q}$, by a theorem of Borel and Harish-Chandra [7, Theorem 7.8] the group $\textrm{Sp}_{2g}(\mathbb{Z})$ is a lattice in $\textrm{Sp}_{2g}(\mathbb{R})$. By Borel density theorem (see [6]) we have that $\textrm{Sp}_{2g}(\mathbb{Z})$ is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{R})$, and so it is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{Q})$. This implies that $W$ is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Q})$. Since the action of $\textrm{Sp}_{2g}(\mathbb{Q})$ on $\mathbb{Q}^{2g}$ is irreducible (see [20, Proposition 3.2]), we get $W=\mathbb{Q}^{2g}$ which concludes the proof.
@@
If the image of the modular map $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$, the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible.
@@
This is an easy consequence of 2.2 and 2.3.
@@
To be more precise, since the group $\textrm{Mon}(\mathcal{A})$ is non-trivial then the functions $\omega_{1},\ldots,\omega_{2g}$ cannot be all defined continuously on the whole of $S(\mathbb{C})$. Moreover, when the action of $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible, neither one of $\omega_{1},\ldots,\omega_{2g}$ nor any single non-zero element of the period lattice can be defined on the whole of $S(\mathbb{C})$. This last thing is also proved in [16, Lemma 5.6] for any abelian scheme over a curve without using irriducibility properties of the monodromy action.
2.4 Relative monodromy of abelian logarithms
Let us consider the exponential map $\exp:\operatorname{Lie}(\mathcal{A})\to\mathcal{A}$ defined in Section 2.2. Notice that it is a topological cover. Thus, it induces a monodromy action of the fundamental group $\pi_{1}(\mathcal{A}(\mathbb{C}),p)$ on the fiber $\exp^{-1}(p)$, where we assume $\phi(p)=s$. Now, let $\sigma:S\to\mathcal{A}$ be a section of the abelian scheme. By restricting the exponential map to $\sigma(S)$ and identifying the fundamental groups $\pi_{1}(\sigma(S(\mathbb{C})),p)$ and $\pi_{1}(S(\mathbb{C}),s)$, we obtain a monodromy action of $\pi_{1}(S(\mathbb{C}),s)$ on the fiber $\exp^{-1}(\sigma(s))$.
Recalling the definition of logarithm given in Equation 5, the set $\exp^{-1}(\sigma(s))$ is the set of all possible determinations of the abelian logarithm of $\sigma$ at the point $s$. Every two such determinations differ each other by an element of the lattice $\Lambda_{s}$, thus we can identify the set $\exp^{-1}(\sigma(s))$ with $\mathbb{Z}^{2g}$. Therefore, we get the monodromy action
@@ 2.4 Relative monodromy of abelian logarithms
which is determined by the analytic continuation of a fixed branch of $\log_{\sigma}(s)$ along loops representing classes in $\pi_{1}(S(\mathbb{C}),s)$.
To be more explicit, fix a determination $\log_{\sigma}(s)$ of the logarithm of $\sigma$ at $s$. The analytic continuation of $\log_{\sigma}(s)$ along a loop representing some $h\in\pi_{1}(S(\mathbb{C}),s)$ is of the form
$\log_{\sigma}+u^{h}_{1}\omega_{1}+\cdots+u^{h}_{2g}\omega_{2g},$ (12)
where $(u^{h}_{1},\ldots,u^{h}_{2g})^{\top}=c(h)$ is the column vector obtained by means of Equation 11.
Now, let us look at the simultaneous monodromy action of $G:=\pi_{1}(S(\mathbb{C}),s)$ on periods and logarithm. By Equation 8 and Equation 11, we can provide a new representation
$\theta_{\sigma}:G\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z}),$
@@
where $\rho(h)\in\textrm{Sp}_{2g}(\mathbb{Z}),c(h)\in\mathbb{Z}^{2g}$. Note that the matrix $\rho(h)$ acts on the periods and does not depend on $\sigma$. Moreover, the vector $c(h)$ encodes the action of $h$ on determinations of the logarithm $\log_{\sigma}$. Define the monodromy group of the section $\sigma$ as $M_{\sigma}:=\theta_{\sigma}(G)$.
Let’s continue considering an abelian scheme $\mathcal{A}\rightarrow S$ and a non-zero section $\sigma:S\rightarrow\mathcal{A}$. Generally neither a basis of periods nor a logarithm $\log_{\sigma}$ can be globally defined on the whole of $S(\mathbb{C})$ (see 2.5 and 2.8 below), but obviously they can be globally defined on the universal cover of $S$. Studying the related monodromy problems corresponds to finding out the minimal (unramified) cover of $S$ on which both a basis of the periods and a logarithm of the section can be defined. With this in mind, we first call $S^{*}\rightarrow S$ the minimal (unramified) cover of $S$ on which a basis for the period lattice can be globally defined and we set $S_{\sigma}\rightarrow S^{*}$ to be the minimal cover of $S^{*}$ where one can define the logarithm of $\sigma$. The tower of covers is represented in the diagram:
$S_{\sigma}\rightarrow S^{*}\rightarrow S.$ (14)
In particular, the group $\textnormal{Mon}(\mathcal{A})$ corresponds to the Galois group of the covering map $S^{*}\rightarrow S$, while the group $M_{\sigma}$ corresponds to the Galois group of the covering map $S_{\sigma}\rightarrow S$. Our interest is in studying the relative monodromy of the logarithm of a section with respect to the monodromy of periods, i.e. studying the covering map $S_{\sigma}\rightarrow S^{*}$. Topologically, this is the same as looking at the variation of logarithm via analytic continuation along loops on $S(\mathbb{C})$ which leave periods unchanged. Moreover, in terms of Equation 8 and Equation 13 this corresponds to studying the group $M_{\sigma}^{\textnormal{rel}}:=\theta_{\sigma}(\ker{\rho})$, which we define as relative monodromy group of $\sigma$.
@@
Note that the Zariski-closures of the discrete groups $\textnormal{Mon}(\mathcal{A})$ and $M_{\sigma}$ as far as the kernel of the projection $\textrm{ker}(\overline{M_{\sigma}}\to\overline{\textnormal{Mon}(\mathcal{A})})$ are the differential Galois groups of some Picard-Vessiot extensions of $\mathbb{C}(S)$ obtained with $\omega_{1},\ldots,\omega_{2g},\log_{\sigma}$ and their derivatives (see [11] for further details about differential Galois theory and Picard-Vessiot extensions). In these terms, the differential Galois group $\textrm{ker}(\overline{M_{\sigma}}\to\overline{\textnormal{Mon}(\mathcal{A})})$ was just studied in [1], [4], [5]; anyway, those results give no information on the relative monodromy of the logarithm over $S^{*}$ because they involve the Zariski closures of $M_{\sigma}$ and $\textnormal{Mon}(\mathcal{A})$: in fact, in [10] the authors pointed out that it may happen that the group $\textrm{ker}(\overline{M_{\sigma}}\to\overline{\textnormal{Mon}(\mathcal{A})})$ contains quite limited information on $M_{\sigma}^{\textrm{rel}}$ since the former may be larger than expected in comparison with the latter. Thus, in order to obtain information on $M_{\sigma}^{\textnormal{rel}}$, we have to introduce considerations of different nature with respect to Bertrand’s and André’s theorems.
@@
Let $\mathcal{A}\to S$ be an abelian scheme. We use Lang-Néron theorem [23, Thorem 1] to show that a non-zero algebraic section $\sigma:S\rightarrow\mathcal{A}$ which is not contained in the fixed part cannot admit a globally defined logarithm on $S(\mathbb{C})$. In fact, assume by contradiction that $\log_{\sigma}$ exists globally on $S(\mathbb{C})$, then $\frac{\sigma}{n}:=\exp\circ\left(\frac{1}{n}\log_{\sigma}\right)$ for any $n\in\mathbb{Z}\setminus\{0\}$ is a holomorphic section such that $n\cdot\frac{\sigma}{n}=\sigma$. This means that $\sigma$ is infinitely divisible in $\Sigma(S)$ contradicting Lang-Néron theorem.
By Equation 12, fixed a determination of $\log_{\sigma}$ on an open set $U$ any element $h\in G:=\pi_{1}(S(\mathbb{C}),s)$ acts by analytic continuation on it in the following way:
$h\cdot\log_{\sigma}=\log_{\sigma}+u_{1}^{h}\omega_{1}+\cdots+u_{2g}^{h}\omega_{2g},$
where $u_{1}^{h},\ldots,u_{2g}^{h}\in\mathbb{Z}$. Choose $h_{1},h_{2}\in G$; by looking at the action of $h_{1}h_{2}$ and recalling Equation 9, we have
Therefore, we obtain the relation
$(u_{1}^{h_{1}h_{2}},\ldots,u_{2g}^{h_{1}h_{2}})=(u_{1}^{h_{1}},\ldots,u_{2g}^{h_{1}})+(u_{1}^{h_{2}},\ldots,u_{2g}^{h_{2}})\rho(h_{1})^{\top}.$
In other words the map $h\mapsto(u_{1}^{h},\ldots,u_{2g}^{h})$ is a cocycle for the described action of $G$ on $\mathbb{Z}^{2g}$. Moreover, in the next proposition we prove that a section admitting a globally defined logarithm is characterized by the fact that the map $h\mapsto(u_{1}^{h},\ldots,u_{2g}^{h})$ is a coboundary for the aforementioned action, i.e. there exists a fixed vector $(u_{1},\ldots,u_{2g})\in\mathbb{Z}^{2g}$ such that $(u_{1}^{h},\ldots,u_{2g}^{h})=(u_{1},\ldots,u_{2g})(\rho(h)^{\top}-1_{2g})$ for all $h\in G$. Thus, the just defined cocycle determines an element in the cohomology group $H^{1}(G,\mathbb{Z}^{2g})$ that describes the obstruction for a(n analytic) section to have a globally defined logarithm.
@@
Suppose that $\sigma$ admits a globally defined logarithm $\ell:S(\mathbb{C})\rightarrow\mathbb{C}^{g}$. Then the two determinations $\ell$ and $\log_{\sigma}$ over $U$ differ by a period $\omega:=n_{1}\omega_{1}+\cdots+n_{2g}\omega_{2g}$, i.e.
$\log_{\sigma}=\ell+n_{1}\omega_{1}+\cdots+n_{2g}\omega_{2g}.$
For $h\in G$, we have
@@ 3.1 Invariance results
In this section we want to provide some preliminary results about the relative monodromy group $M_{\sigma}^{\textnormal{rel}}$ in order to prove that our main theorem are insensitive to some natural operations. First of all, let us notice that 1.2 is invariant under isogenies: this was proven in [33, Theorem 3.6]. In what follows we describe other invariance results.
@@
$\widetilde{\omega_{1}}=\omega_{1}\circ\varphi,\quad\ldots,\quad\widetilde{\omega_{2g}}=\omega_{2g}\circ\varphi.$ (15)
@@
Let $\sigma:S\rightarrow\mathcal{A}$ be a non-torsion section of $\mathcal{A}\rightarrow S$. Since $\widetilde{\mathcal{A}}\rightarrow\widetilde{S}$ is obtained as pullback of $\mathcal{A}\rightarrow S$, then the abelian varieties $\widetilde{\mathcal{A}}_{\widetilde{s}}$ and $\mathcal{A}_{s}$ are canonically identified; therefore the pullback $\varphi^{*}(\sigma):=\sigma\circ\varphi:\widetilde{S}\rightarrow\widetilde{\mathcal{A}}$ is a non-torsion section of $\widetilde{\mathcal{A}}\rightarrow\widetilde{S}$. We have the associated monodromy representations
$\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z}),\qquad\theta_{\varphi^{*}(\sigma)}:\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s})\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z}).$
Using again the fact that the abelian varieties $\widetilde{\mathcal{A}}_{\widetilde{s}}$ and $\mathcal{A}_{s}$ are canonically identified, we obtain that a determination of the logarithm of $\varphi^{*}(\sigma)$ at $\widetilde{s}$ can be defined by the equation
$\log_{\varphi^{*}(\sigma)}(\widetilde{s}):=\log_{\sigma}(s).$ (17)
Hence, we have $\theta_{\varphi^{*}(\sigma)}(h)=\theta_{\sigma}(\varphi_{*}(h))$ for each $h\in\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s})$. Now, let us consider the relative monodromy group
$M_{\varphi^{*}(\sigma)}^{\textnormal{rel}}=\theta_{\varphi^{*}(\sigma)}(\ker{\widetilde{\rho}}).$
@@
$M_{\varphi^{*}(\sigma)}^{\textnormal{rel}}=\theta_{\sigma}(\varphi_{*}(\ker{\widetilde{\rho}}))=\theta_{\sigma}(\ker{\rho}\cap\varphi_{*}(\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s}))).$
Since $\ker{\rho}\cap\varphi_{*}(\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s}))\subset\ker{\rho}$, we get $M_{\varphi^{*}(\sigma)}^{\textnormal{rel}}\subseteq M_{\sigma}^{\textnormal{rel}}$. The conclusion follows.
@@
$(ii)$: Fixed a branch $\ell_{\sigma}$ of $\log_{\sigma}$, a determination $\ell_{n\sigma}$ of the logarithm of $n\sigma$ can be defined by the equation
$\ell_{n\sigma}=n\ell_{\sigma}.$
For any loop $\alpha$ representing a homotopy class $h\in\pi_{1}(S(\mathbb{C}),s)$ we have the corresponding variations:

### v4

@@ 3.2 The case of universal families
Recall that the monodromy group of the universal family $\textrm{Mon}(\mathfrak{A}_{g})$ is a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$. When the modular map $p:S\to\mathbb{A}_{g}$ is finite, the monodromy group $\textrm{Mon}(\mathcal{A})$ of $\mathcal{A}\to S$ is still a finite index subgroup of $\textrm{Sp}_{2g}(\mathbb{Z})$ thanks to Equation 12; in particular, $\textrm{Mon}(\mathcal{A})$ is Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$. Moreover, when the image of the modular map $T$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$, the monodromy group $\textrm{Mon}(\mathcal{A})$ is still Zariski-dense in $\textrm{Sp}_{2g}(\mathbb{Z})$ by [4, Remark 10.1.2].
@@
Assume that there exists a non-trivial subvector space $W$ of $\mathbb{Q}^{2g}$ which is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Z})$. Since the symplectic group $\textrm{Sp}_{2g}$ is a simple algebraic Lie group defined over $\mathbb{Q}$, by a theorem of Borel and Harish-Chandra [8, Theorem 7.8] the group $\textrm{Sp}_{2g}(\mathbb{Z})$ is a lattice in $\textrm{Sp}_{2g}(\mathbb{R})$. By Borel density theorem (see [7]) we have that $\textrm{Sp}_{2g}(\mathbb{Z})$ is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{R})$, and so it is Zariski dense in $\textrm{Sp}_{2g}(\mathbb{Q})$. This implies that $W$ is fixed by the action of $\textrm{Sp}_{2g}(\mathbb{Q})$. Since the action of $\textrm{Sp}_{2g}(\mathbb{Q})$ on $\mathbb{Q}^{2g}$ is irreducible (see [27, Proposition 3.2]), we get $W=\mathbb{Q}^{2g}$ which concludes the proof.
@@
If the image of the modular map $T$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$, the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible.
@@
This is an easy consequence of 3.2 and 3.3.
@@
To be more precise, since the group $\textrm{Mon}(\mathcal{A})$ is non-trivial then the functions $\omega_{1},\ldots,\omega_{2g}$ cannot be all defined continuously on the whole of $S(\mathbb{C})$. Moreover, when the action of $\textrm{Mon}(\mathcal{A})$ on the lattice of periods is irreducible, neither one of $\omega_{1},\ldots,\omega_{2g}$ nor any single non-zero element of the period lattice can be defined on the whole of $S(\mathbb{C})$. This last thing is also proved in [21, Lemma 5.6] for any abelian scheme over a curve without using irriducibility properties of the monodromy action.
3.3 Monodromy groups of sections
Let us consider the exponential map $\exp:\operatorname{Lie}(\mathcal{A})\to\mathcal{A}(\mathbb{C})$ defined in Section 2.2. Notice that it is a topological cover. Thus, it induces a monodromy action of the fundamental group $\pi_{1}(\mathcal{A}(\mathbb{C}),p)$ on the fiber $\exp^{-1}(p)$, where we assume $\phi(p)=s$. Let $\sigma:S\to\mathcal{A}$ be a section of the abelian scheme, we obtain a monodromy action of $\pi_{1}(S(\mathbb{C}),s)$ on the fiber $\exp^{-1}(\sigma(s))$.
Recalling the definition of logarithm given in Equation 7, the set $\exp^{-1}(\sigma(s))$ is the set of all possible determinations of the abelian logarithm of $\sigma$ at the point $s$. Every two such determinations differ each other by an element of the lattice $\Lambda_{s}$, therefore, after fixing a branch $\log_{\sigma}(s)\in\exp^{-1}(\sigma(s))$ of the logarithm at $s$, we get the monodromy action
$\rho^{\prime}:\pi_{1}(S(\mathbb{C}),s)\to\operatorname{Aut}(\mathbb{Z}^{2g})\,.$
The representation $\rho^{\prime}$ induces a map (which is not a group homomorphism):
@@ 3.3 Monodromy groups of sections
$h$ $\mapsto\rho^{\prime}(h)(\log_{\sigma}(s))\,.$
The function $c$ can be described explicitly. In fact, any element $h\in\pi_{1}(S(\mathbb{C}),s)$ acts by analytic continuation on $\log_{\sigma}(s)$ in the following way:
$h\cdot\log_{\sigma}=\log_{\sigma}+u_{1}^{h}\omega_{1}+\cdots+u_{2g}^{h}\omega_{2g},$ (14)
where $u_{1}^{h},\ldots,u_{2g}^{h}\in\mathbb{Z}$. Then $c(h)=(u^{h}_{1},\ldots,u^{h}_{2g})^{\top}$. Again, since $S(\mathbb{C})$ is path connected we can fix the base point $s\in S(\mathbb{C})$ once and for all.
Moreover, if we choose $h_{1},h_{2}\in\pi_{1}(S(\mathbb{C}),s)$ we can consider the action of the product $h_{1}h_{2}$. Recalling Equation 11, we have
Therefore, we obtain the cocycle relation
$(u_{1}^{h_{1}h_{2}},\ldots,u_{2g}^{h_{1}h_{2}})=(u_{1}^{h_{1}},\ldots,u_{2g}^{h_{1}})+(u_{1}^{h_{2}},\ldots,u_{2g}^{h_{2}})\rho(h_{1})^{\top}.$ (15)
Now, let us look at the simultaneous monodromy action of $\pi_{1}(S(\mathbb{C}),s)$ on periods and logarithm. Thanks to Equation 15, using the functions described in Equation 10 and Equation 13, we can provide a new representation
$\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z}),$
@@
for $\rho(h)\in\textrm{Sp}_{2g}(\mathbb{Z}),c(h)\in\mathbb{Z}^{2g}$. Note that the matrix $\rho(h)$ acts on the periods and does not depend on $\sigma$. Moreover, the vector $c(h)$ encodes the action of $h$ on determinations of the logarithm $\log_{\sigma}$.
Definition 3.6.
The monodromy group of the section $\sigma$ is $M_{\sigma}:=\theta_{\sigma}(\pi_{1}(S(\mathbb{C}),s))$.
Let’s continue considering an abelian scheme $\mathcal{A}\rightarrow S$ and a non-zero section $\sigma:S\rightarrow\mathcal{A}$. Generally neither a basis of periods nor a logarithm $\log_{\sigma}$ can be globally defined on the whole of $S(\mathbb{C})$ (see 3.5 and 3.10 below), but obviously they can be globally defined on the universal cover of $S(\mathbb{C})$. Studying the related monodromy problems corresponds to finding out the minimal topological cover of $S(\mathbb{C})$ on which both a basis of the periods and a logarithm of the section can be defined. With this in mind, we first call $S^{*}\rightarrow S(\mathbb{C})$ the minimal topological cover of $S(\mathbb{C})$ on which a basis for the period lattice can be globally defined and we set $S_{\sigma}\rightarrow S^{*}$ to be the minimal cover of $S^{*}$ where one can define the logarithm of $\sigma$. The tower of covers is represented in the diagram:
$S_{\sigma}\rightarrow S^{*}\rightarrow S(\mathbb{C}).$ (17)
In particular, the group $\textnormal{Mon}(\mathcal{A})$ corresponds to the Galois group of the covering map $S^{*}\rightarrow S(\mathbb{C})$, while the group $M_{\sigma}$ corresponds to the Galois group of the covering map $S_{\sigma}\rightarrow S(\mathbb{C})$. Our interest is in studying the relative monodromy of the logarithm of a section with respect to the monodromy of periods, i.e. studying the covering map $S_{\sigma}\rightarrow S^{*}$. Topologically, this is the same as looking at the variation of logarithm via analytic continuation along loops on $S(\mathbb{C})$ which leave periods unchanged. With the notations of Equation 10 and Equation 16 this corresponds to studying group defined below:
Definition 3.7.
$M_{\sigma}^{\textnormal{rel}}:=\theta_{\sigma}(\ker{\rho})$ is called the relative monodromy group of $\sigma$.
@@
Note that the Zariski-closures of the discrete groups $\textnormal{Mon}(\mathcal{A})$ and $M_{\sigma}$ as far as the kernel of the projection $\textrm{ker}(\overline{M_{\sigma}}\to\overline{\textnormal{Mon}(\mathcal{A})})$ are the differential Galois groups of some Picard-Vessiot extensions of $\mathbb{C}(S)$ obtained with $\omega_{1},\ldots,\omega_{2g},\log_{\sigma}$ and their derivatives (see [14] for further details about differential Galois theory and Picard-Vessiot extensions). In these terms, the differential Galois group $\textrm{ker}(\overline{M_{\sigma}}\to\overline{\textnormal{Mon}(\mathcal{A})})$ was just studied in [2], [5], [6]. Anyway, those results give no information on the relative monodromy of the logarithm over $S^{*}$ because they involve the Zariski closures of $M_{\sigma}$ and $\textnormal{Mon}(\mathcal{A})$: in fact, in [13] the authors pointed out that it may happen that the group $\textrm{ker}(\overline{M_{\sigma}}\to\overline{\textnormal{Mon}(\mathcal{A})})$ contains quite limited information on $M_{\sigma}^{\textrm{rel}}$ since the former may be larger than expected in comparison with the latter. Thus, in order to obtain information on $M_{\sigma}^{\textnormal{rel}}$, we have to introduce considerations of different nature with respect to Bertrand’s and André’s theorems.
@@
Let $\mathcal{A}\to S$ be an abelian scheme. We use Lang-Néron theorem [30, Thorem 1] to show that a non-zero algebraic section $\sigma:S\rightarrow\mathcal{A}$ which is not contained in the fixed part cannot admit a globally defined logarithm on $S(\mathbb{C})$. In fact, assume by contradiction that $\log_{\sigma}$ exists globally on $S(\mathbb{C})$, then $\frac{\sigma}{n}:=\exp\circ\left(\frac{1}{n}\log_{\sigma}\right)$ for any $n\in\mathbb{Z}\setminus\{0\}$ is a holomorphic section such that $n\cdot\frac{\sigma}{n}=\sigma$. This means that $\sigma$ is infinitely divisible in $\Sigma(S)$ contradicting Lang-Néron theorem.
Fix a determination of $\log_{\sigma}$ on an open set $U$ containing the base point $s$ and use the notation $G:=\pi_{1}(S(\mathbb{C}),s)$. Thanks to Equation 15, the map $c:h\mapsto(u_{1}^{h},\ldots,u_{2g}^{h})$ is a cocycle for the action of $G$ on $\mathbb{Z}^{2g}$ described in Equation 14. Moreover, in the next proposition we prove that a section admitting a globally defined logarithm is characterized by the fact that the map $h\mapsto(u_{1}^{h},\ldots,u_{2g}^{h})$ is a coboundary for the aforementioned action, i.e. there exists a fixed vector $(u_{1},\ldots,u_{2g})\in\mathbb{Z}^{2g}$ such that $(u_{1}^{h},\ldots,u_{2g}^{h})=(u_{1},\ldots,u_{2g})(\rho(h)^{\top}-1_{2g})$ for all $h\in G$. Thus, the just defined cocycle determines an element in the cohomology group $H^{1}(G,\mathbb{Z}^{2g})$ that describes the obstruction for a(n analytic) section to have a globally defined logarithm.
@@
Suppose that $\sigma$ admits a globally defined logarithm $\ell:S(\mathbb{C})\rightarrow\mathbb{C}^{g}$. Then the two determinations $\ell$ and $\log_{\sigma}$ over $U$ differ by a period $\omega:=n_{1}\omega_{1}+\cdots+n_{2g}\omega_{2g}$, i.e. $\log_{\sigma}=\ell+n_{1}\omega_{1}+\cdots+n_{2g}\omega_{2g}$. For $h\in G$, we have
@@ 4.1 Invariance results
In this section we want to provide some preliminary results about the relative monodromy group $M_{\sigma}^{\textnormal{rel}}$ in order to prove that our main theorem are insensitive to some natural operations. First of all, let us notice that 1.2 is invariant under isogenies: this was proven in [41, Theorem 3.6]. In what follows we describe other invariance results.
@@
$\widetilde{\omega_{1}}=\omega_{1}\circ\varphi,\;\ldots,\;\widetilde{\omega_{2g}}=\omega_{2g}\circ\varphi.$ (18)
@@
Let $\sigma:S\rightarrow\mathcal{A}$ be a non-torsion section of $\mathcal{A}\rightarrow S$. Since $\widetilde{\mathcal{A}}\rightarrow\widetilde{S}$ is obtained as pullback of $\mathcal{A}\rightarrow S$, then the abelian varieties $\widetilde{\mathcal{A}}_{\widetilde{s}}$ and $\mathcal{A}_{s}$ are canonically identified; therefore the pullback $\varphi^{*}\sigma:\widetilde{S}\rightarrow\widetilde{\mathcal{A}}$ is a non-torsion section of $\widetilde{\mathcal{A}}\rightarrow\widetilde{S}$. We have the associated monodromy representations
$\theta_{\sigma}:\pi_{1}(S(\mathbb{C}),s)\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z}),\qquad\theta_{\varphi^{*}\sigma}:\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s})\rightarrow\textrm{SL}_{2g+1}(\mathbb{Z}).$
Using again the fact that the abelian varieties $\widetilde{\mathcal{A}}_{\widetilde{s}}$ and $\mathcal{A}_{s}$ are canonically identified, we obtain that a determination of the logarithm of $\varphi^{*}\sigma$ at $\widetilde{s}$ can be defined by the equation
$\log_{\varphi^{*}\sigma}(\widetilde{s}):=\log_{\sigma}(s).$ (20)
Hence, we have $\theta_{\varphi^{*}\sigma}(h)=\theta_{\sigma}(\varphi_{*}(h))$ for each $h\in\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s})$. Now, let us consider the relative monodromy group
$M_{\varphi^{*}\sigma}^{\textnormal{rel}}=\theta_{\varphi^{*}\sigma}(\ker{\widetilde{\rho}}).$
@@
$M_{\varphi^{*}\sigma}^{\textnormal{rel}}=\theta_{\sigma}(\varphi_{*}(\ker{\widetilde{\rho}}))=\theta_{\sigma}(\ker{\rho}\cap\varphi_{*}(\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s}))).$
Since $\ker{\rho}\cap\varphi_{*}(\pi_{1}(\widetilde{S}(\mathbb{C}),\widetilde{s}))\subset\ker{\rho}$, we get $M_{\varphi^{*}\sigma}^{\textnormal{rel}}\subseteq M_{\sigma}^{\textnormal{rel}}$. The conclusion follows.
@@
$(ii)$: Fixed a branch $\ell_{\sigma}$ of $\log_{\sigma}$, a determination $\ell_{n\sigma}$ of the logarithm of $n\sigma$ can be defined by the equation $\ell_{n\sigma}=n\ell_{\sigma}$. For any loop $\alpha$ representing a homotopy class $h\in\pi_{1}(S(\mathbb{C}),s)$ we have the corresponding variations: